OpenStudy (anonymous):

Can someone please help me?? Will Fan AND Medal!!! Which functions have a positive rate of change and which have a negative rate of change? Column A Column B 1. 2x − 5y = −15 2. y = −4x + 9 3. x y 2 –10 4 –16 6 –22 4. x y –3 15 –2 11 –1 7 A. positive B. negative

OpenStudy (xapproachesinfinity):

hint: Rate of change is the slope

OpenStudy (anonymous):

how do I find it?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

if i give this kind of equation y=2x-1 what is the slope?

OpenStudy (anonymous):

2?

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (anonymous):

so then it is positive?

OpenStudy (xapproachesinfinity):

2x-5y=15 solve for y to have the same form that one have

OpenStudy (anonymous):

how?

OpenStudy (xapproachesinfinity):

well of course it is positive do you see any negaative sign there?

OpenStudy (anonymous):

I was talking about the equation.

OpenStudy (xapproachesinfinity):

y=-2x+1 this has negative slope

OpenStudy (anonymous):

oh

OpenStudy (xapproachesinfinity):

well how do you solve a linear equation

OpenStudy (anonymous):

?

OpenStudy (xapproachesinfinity):

2x-5y=-15 what do you do to solve for y?

OpenStudy (xapproachesinfinity):

you need to solve for y to find the slope we want it to be of this form y=mx+b where m=slope and b is y intercept

OpenStudy (anonymous):

oh I got it.

OpenStudy (xapproachesinfinity):

ok!

OpenStudy (xapproachesinfinity):

you gonna do the same for the other ones

OpenStudy (xapproachesinfinity):

is that a table with two columns?

OpenStudy (xapproachesinfinity):

the way you wrote it is vague

OpenStudy (xapproachesinfinity):

if you have trouble making a table take a picture of your problem and post it